### v2

Abstract
Current state-of-the-art segmentation models encode entire images before focusing on specific objects. As a result, they waste computational resources—particularly when small objects are to be segmented in high-resolution scenes. We introduce FLIP (Fovea-Like Input Patching), a parameter-efficient vision model that realizes object segmentation through biologically-inspired top-down attention. FLIP selectively samples multi-resolution patches centered on objects of interest from the input. As a result, it allocates high-resolution processing to object centers while maintaining coarser peripheral context. This off-grid, scale-invariant design enables FLIP to outperform META’s Segment Anything models (SAM) by large margins: With more than 1000x fewer parameters, FLIP-Tiny (0.51M parameters) reaches a mean IoU of 78.24% while SAM-H reaches 75.41% IoU (641.1M parameters). FLIP-Large even achieves 80.33% mean IoU (96.6M parameters), still running about 6$\times$ faster than SAM-H. We evaluate on six benchmarks in total. In five established benchmarks (Hypersim, KITTI-360, OpenImages, COCO, LVIS) FLIP consistently outperforms SAM and various variants of it. In our novel ObjaScale dataset, which stress-tests scale invariance with objects ranging from 0.0001% up-to 25% of the image area, we show that FLIP segments even very small objects accurately, where existing models fail severely. FLIP opens new possibilities for real-time, object-centric vision applications and offers much higher energy efficiency. We believe that FLIP can act as a powerful foundation model, as it is very well-suited to track objects over time, for example, when being integrated into slot-based scene segmentation architectures.
1 Introduction
Object-centric models have emerged as a powerful paradigm for structured perception in visual tasks. They offer the potential to represent complex scenes in a more interpretable and compositional manner. While traditional architectures such as Convolutional Neural Networks (CNNs) Liu et al. (2022) and Vision Transformers (ViTs) Dosovitskiy et al. (2020) have demonstrated impressive performance on large-scale datasets, they often lack the nuanced object-level understanding required for robust scene parsing. Furthermore, these models typically require massive amounts of labeled data and exhibit vulnerabilities to adversarial perturbations.
Figure 1: Comparison of mean IoU across six diverse datasets (Hypersim, KITTI-360, OpenImages, COCO, LVIS, ObjaScale) plotted against model size (in millions of parameters). FLIP variants consistently outperform SAM variants while using several orders of magnitude fewer parameters.
Recent advances in object-centric learning, such as Slot Attention Locatello et al. (2020), have sought to address these challenges by enabling the model to discover and represent objects within a scene as distinct entities. Models like SAVi++ Elsayed et al. (2022), VideoSAUR Zadaianchuk et al. (2024), and Loci Traub et al. (2023; 2024a; 2024b) have advanced the state-of-the-art in unsupervised object-centric learning. They are able to disentangle objects from complex backgrounds and track their identities over time. Still, these approaches struggle to scale effectively to more complex, real-world data.
In contrast, the Segment Anything (SAM) model Kirillov et al. (2023) has introduced a paradigm shift in object-centric learning. SAM learns from a vast array of diverse data that is segmented by a two stage segmentation process. First, a powerful transformer-based foundational model encodes the complete image. Second, a query-based focusing mechanism specifies which object to segment. Only this second mechanism targets one image area or object and leads to the production of the targeted output mask. SAM marks the state-of-the-art in object segmentation tasks. However, despite its impressive performance, SAM has its limits. First, the transformer-based encoder requires very large computational resources. Second, the encoder backbone encodes the complete image, potentially wasting processing resources, particularly when small objects are to be segmented.
Figure 2: Visualization of our FLIP (Fovea-Like Input Patching) approach applied to an image from the KITTI-360 dataset, showcasing potential applications in autonomous driving. The figure illustrates how our model dynamically focuses on multiple objects within a complex urban scene by allocating multi-resolution patches centered around estimated object locations. Higher-resolution patches (smaller sizes) are concentrated on critical areas such as vehicles and road-signs, emulating a foveal vision system, while lower-resolution patches (larger sizes) cover peripheral regions to enable the consideration of the surrounding context. Patches are color-coded by size: purple for $16\times 16$ patches, yellow for $8\times 8$, green for $4\times 4$, blue for $2\times 2$, and red for $1\times 1$.
Derivative model variants like EfficientSAM, MobileSAM, and FastSAM Xiong et al. (2023); Zhang et al. (2023); Zhao et al. (2023) address the former limitation. In our evaluation of single-object segmentation performance across multiple datasets, EfficientSAM employs SAMI (SAM-leveraged masked image pretraining), training lightweight ViT encoders to reconstruct features from SAM’s ViT-H encoder, and achieves 72.29% mean IoU with only 10.22M parameters (EfficientSAM-T) compared to SAM-H’s 75.41% IoU with 641M parameters. MobileSAM applies decoupled distillation, replacing SAM’s ViT-H encoder (632M parameters) with TinyViT (5M parameters) Wu et al. (2022), achieving 71.33% mean IoU with an over 60$\times$ parameter reduction and $\sim$21ms runtime per image. FastSAM replaces the first-stage transformer-based encoder in SAM with a convolutional ANN (CNN) that is pre-trained to segment full images. It achieves a runtime per image of under 10ms but exhibits significantly reduced performance with only 44.58% mean IoU (FastSAM-s).
At this point nearly all segmentation techniques, including SAM and its efficient variants, rely on a full image encoder. This is also the case for most object centric models, such as SlotAttention, VideoSAUR, and related work Locatello et al. (2020); Elsayed et al. (2022); Singh et al. (2022); Zadaianchuk et al. (2024), which first encode the entire image before assigning information to slots. The challenge to computationally efficiently segment and track small but potentially high-resolution objects across diverse and complex scenes remains. Here, we introduce a vision transformer-based foundational model that utilizes a local, pixel-patch-based image encoder to generate accurate masks.
Figure 3: Overview of the FLIP architecture. The Foveal Patching module dynamically samples multi-resolution patches centered around objects of interest. These patches are embedded into a unified latent space using resolution-specific Patch Embedding Modules ($E_{r_{0}})$ to $E_{r_{K}}$). The Vision Transformer Encoder processes the embedded patches, generating keys $K_{1..n}$ and values $V_{1..n}$. The Pixel-Predictor performs attention over queries derived from pixel coordinates $Q_{x,y}$ enabling instance segmentation with pixel-level precision.
In particular, we introduce FLIP: a fovea-like input patching approach that is integrated in an object-centric, off-grid vision framework. FLIP dynamically adapts its processing pipeline to the object’s size and spatial characteristics. It ignores currently irrelevant image subregions and focuses on critical regions with a flexible, multi-resolution approach. Our key contributions are:
  • Off-Grid, Scale-Invariant Object Encoding: We introduce a fovea-inspired pixel-based patch sampling method that directly encodes image regions off-grid, adaptively focusing on objects of interest in a multi-resolution fashion. This scale-invariant approach is robust to large variations in object size and image resolution. It enables the detailed encoding of very small objects even in high-resolution scenes.
  • Exceptional Parameter and Computational Efficiency: FLIP achieves superior segmentation performance while using orders of magnitude fewer parameters than existing state-of-the-art models. With lightweight encoders and efficient processing, FLIP provides significant energy savings and faster inference times, making it suitable for real-time applications and resource-constrained environments.
  • State-of-the-Art Segmentation Performance with High Parameter Efficiency: Despite using significantly fewer parameters compared to state-of-the-art models from the SAM family, FLIP achieves superior segmentation accuracy on standard benchmarks such as Hypersim, KITTI-360, OpenImages, COCO and LVIS.
  • Superior performance on novel dataset ObjaScale We introduce ObjaScale, a high-resolution evaluation benchmark with Blender-rendered objects at varying scales on real-world High Dynamic Range Image (HDRI) backgrounds, designed to stress-test scale invariance. As in the other datasets, FLIP outperforms all SAM variants on ObjaScale and shows that it can segment even rather tiny objects accurately.
2 Related Work
Several lines of research contain methods that share thematically similar ideas in handling multi-scale objects, dynamic sampling, or biologically inspired foveation. Deformable Convolutional Networks Dai et al. (2017) and their successors Zhu et al. (2019); Xiong et al. (2024) introduce learnable offsets or selective kernels to better handle varying spatial structures. Focal Sparse Convolutional Networks Chen et al. (2022) focus computation on salient 3D regions in point clouds. Additionally, biologically inspired foveation has been explored in works like Lukanov et al. (2021); Kaplanyan et al. (2019); Thavamani et al. (2021).
3 Methods
In this section, we present the overall Fovea-Like Input Patching (FLIP) architecture. Its main processing pipeline is shown in Figure 3. FLIP is a supervised vision model for efficient object-centric segmentation. It combines a fovea-inspired patching mechanism with a Vision Transformer (ViT)-based encoder. The encoded information is then used to generate the targeted object’s segmentation mask.
The fovea-like patching mechanism dynamically selects multi-resolution patches based on a 2D Gaussian distribution, which has an effect similar to the query in SAM variants but acts directly on the input image. High-resolution patches focus on the object center, capturing fine details, while coarser patches cover peripheral regions, which inform FLIP about the surrounding context.
The ViT encoder processes the sampled patches and outputs latent tokens from which we compute keys $K_{1..n}$ and values $V_{1..n}$. The Pixel-Predictor then computes queries $Q_{x,y}$ from pixel coordinates and uses an attention mechanism to predict the segmentation mask with pixel level accuracy (see Figure 5).
FLIP is trained end-to-end with a sparse supervised loss for mask prediction. In the following subsections, we provide the mathematical details of the fovea-like sampling mechanism, the ViT encoder, and the Pixel-Predictor.
3.1 Fovea-like Input Patching
We equip our ViT-based model with a multi-resolution, fovea-inspired patching mechanism that centers around the object of interest. The mechanism preserves high-resolution detail at the object center and coarser coverage in peripheral regions. Specifically, we derive the object center $\mu=(\mu_{x},\mu_{y})$ and covariance $\Sigma$ from the ground-truth mask, yielding a 2D Gaussian $\mathcal{N}(\mu,\Sigma)$ that approximates the object’s spatial extent and orientation in compressed form. For input sampling, the Gaussian serves as a spatial input query—similar to the prompts in SAM variants—from which we then draw patches at multiple scales.
We define $K$ patch sizes $p_{1}<p_{2}<\dots<p_{K}$, from smallest (highest resolution) to largest. We fix the total number of patches $N$ and choose $N_{i}$ patches for patch size $p_{i}$ as follows. First, we compute default numbers $\hat{N}_{i}$ by approximating the integral of the 2D Gaussian and dividing by $p_{i}^{2}$, scaled by a coverage parameter $c\in(0.1,2)$ that controls the density of patch sampling. Then, starting at the coarsest resolution (i.e., $p_{K}$), we compute $N_{i}=\min(\hat{N}_{i},N-\sum_{j=i+1}^{K}N_{j})$. As a result, the chosen patches cover the Gaussian inner area with a density proportional to $c$ and the number of chosen patches per size distributes itself from coarse to fine until the total number of patches $N$ is reached ($N=\sum_{i}^{K}N_{i}$).
We sample $N_{i}$ patches from $\mathcal{N}(\mu,\Sigma)$ with continuous coordinates, allowing patches to be centered at real-valued positions $(x,y)\in\mathbb{R}^{2}$. Pixel values are extracted using bilinear interpolation, enabling sub-pixel precision. To control patch overlap, we employ a spatial hash grid that tracks placed patches. When sampling, we reject candidate positions that would result in overlap exceeding the maximum overlap threshold $\tau\in[0,4]$. This ensures efficient spatial coverage while allowing for controlled redundancy. We implement this sampling mechanism as a Python C extension in the data loading pipeline, achieving approximately 1ms average extraction time per image. Each sampled patch is first flattened and then mapped to a common embedding space via resolution-specific encoders $E_{r_{i}}$ (similarly to Jaegle et al., 2021), yielding a set of tokens $T=\{t_{1},...,t_{N}\}$ of equal tensor size. These embedded tokens are concatenated and fed to the main ViT layers. Note that the approach is in principle independent of the full image size, because sampling depends on $\mu$ and $\Sigma$ only.
3.2 Encoder
For encoding the multi-resolution patches, we employ a Vision Transformer architecture with modified attention blocks inspired by LaPE Yu et al. (2023). Our design differs from standard transformers by incorporating position-aware attention without position embedding normalization, maintaining separate layer normalization for queries, keys, and values.
Figure 4 illustrates the architecture of our attention blocks. Each block receives the set of tokens $X_{l}$ from the previous layer as input. The first layer receives $X_{1}=T$, which are the tokens produced by patch-size-specific embedding MLPs $E_{r_{0}}..E_{r_{K}}$ that process the foveal patches. Similar to LaPE, we inject relative positional embeddings $\mathrm{PE}$ into every layer.
Figure 4: Architecture of the Attention block in FLIP. Positional embeddings influence queries, keys and values independent of feature normalization.
The relative positional embeddings $(\hat{x}_{i},\hat{y}_{i}$) are computed with respect to the 2D Gaussian input query. From $\Sigma$, we extract rotation $(\theta_{a},\theta_{b})$ and scale $(\sigma_{x},\sigma_{y})$ via a standard eigenvalue decomposition of the covariance matrix. For each token at position $(x_{i},y_{i})$, we shift coordinates by $(\mu_{x},\mu_{y})$, rotate by $(\theta_{a},\theta_{b})$, and scale by $(\sigma_{x},\sigma_{y})$. We then use separate MLPs to compute positional embeddings for queries, keys and values:
$\mathrm{PE}_{i}^{q,k,v}$ $=\text{MLP}_{i}^{q,k,v}(\hat{x}_{i},\hat{y}_{i}),$
$\mathbf{Q}^{\prime}_{i}$ $=\mathbf{Q}_{i}+\mathrm{PE}_{i}^{q},$
$\mathbf{K}^{\prime}_{i}$ $=\mathbf{K}_{i}+\mathrm{PE}_{i}^{k},$
$\mathbf{V}^{\prime}_{i}$ $=\mathbf{V}_{i}+\mathrm{PE}_{i}^{v},$
where MLP is a two-layer MLP with SiLU activations that maps 2D coordinates to positional embeddings.
Unlike standard transformers, we apply individual layer normalization to queries, keys, and values before projection, which empirically improved performance over a single layer normalization (see Table 2):
$Q$ $=\text{LN}_{Q}(X_{l})W_{Q},$
$K$ $=\text{LN}_{K}(X_{l})W_{K},$
$V$ $=\text{LN}_{V}(X_{l})W_{V},$
where $\text{LN}_{Q}$, $\text{LN}_{K}$, and $\text{LN}_{V}$ are separate LayerNorm modules.
The attention computation proceeds as:
$Q^{\prime}=Q+\mathrm{PE}_{Q},\quad K^{\prime}$ $=K+\mathrm{PE}_{K},\quad V^{\prime}=V+\mathrm{PE}_{V},$
$\text{Attn}(Q^{\prime},K^{\prime},V)$ $=\text{softmax}\left(\frac{Q^{\prime}K^{\prime T}}{\sqrt{d_{k}}}\right)V^{\prime},$
$Y$ $=X_{l}+\alpha\cdot\text{Attn}(Q^{\prime},K^{\prime},V^{\prime})W_{out},$
where $\alpha$ is a learnable scaling parameter initialized to 0.1 for stable training, and $d_{k}$ is the key dimension.
Each attention block is followed by a residual MLP:
$X_{l+1}=Y+\beta\cdot\text{MLP}(\text{LN}(Y)),$
where $\beta$ is another learnable scaling parameter. The MLP uses a bottleneck design with SiLU activations and expansion factor of 4. Depending on the model configuration, we stack 3, 5, 8, or 24 of these blocks to form the complete encoder.
Figure 5: Architecture of the Pixel-Predictor. Encoder features $X_{E}$ are processed through a residual MLP. The attention mechanism computes keys (K) and values (V) enhanced with patch positional embeddings ($\mathrm{PE}_{i}$), while queries (Q) use positional embeddings derived from the target pixels coordinates ($\mathrm{PE}_{t}$). The attended features pass through a residual MLP and a final linear layer to produce mask predictions $M_{t}$ for each query pixel.
3.3 Pixel-Level Mask Prediction
The Pixel-Predictor generates precise segmentation masks by computing attention between encoded patch features and query pixels. Given the encoder output tokens $X_{e}$, we predict mask values for arbitrary pixel coordinates through a sparse attention mechanism.
The predictor consists of preprocessing and postprocessing residual MLP layers surrounding a specialized attention module. We first preprocess the encoder features:
$\hat{X}_{e}=\text{MLP}_{\text{pre}}(X_{e})$
For each target pixel at coordinates $(x_{t},y_{t})$, we compute relative positional embeddings with respect to the input 2D Gaussian by shifting coordinates by $(\mu_{x},\mu_{y})$, rotating by $(\theta_{a},\theta_{b})$, and scaling by $(\sigma_{x},\sigma_{y})$ to get $(\hat{x_{t}},\hat{y_{t}})$:
$\mathrm{PE}_{t}$ $=\text{MLP}_{t}(\hat{x_{t}},\hat{y_{t}}),$
$\mathrm{PE}_{i}$ $=\text{MLP}_{i}(\hat{x}_{i},\hat{y}_{i}),$
where MLP is a two-layer MLP with SiLU activations that maps 2D coordinates to positional embeddings, and $\mathrm{PE}_{i}$ represents the positional embeddings for each encoded and preprocessed input patch ($\hat{X}_{e}$).
In contrast to the encoder, the attention mechanism in the Pixel-Predictor is not residual. We compute embedding-informed keys and values as follows:
$K_{\text{enhanced}}$ $=\text{MLP}_{K}(\text{concat}[\text{LN}_{K}(\hat{X}_{e}),\mathrm{PE}_{i}]),$
$V_{\text{enhanced}}$ $=\text{LN}_{V}(\hat{X}_{e})W_{V}+\text{MLP}_{V}(\mathrm{PE}_{i})$
For each query pixel, we compute:
$Q_{\text{pixel}}$ $=\mathrm{PE}_{t},$
$\text{feat}_{\text{pixel}}$ $=\text{Attention}(Q_{\text{pixel}},K_{\text{enhanced}},V_{\text{enhanced}}),$
$M_{x,y}$ $=\text{Linear}(\text{feat}_{\text{pixel}}+\text{MLP}_{\text{post}}(\text{feat}_{\text{pixel}}))$
where $M_{x,y}$ is the predicted mask logit at position $(x,y)$.
Figure 6: Examples of multi-resolution patch inputs (top row) and corresponding mask predictions (bottom row) from FLIP. Input patches are color-coded by size: purple ($16\times 16$), yellow ($8\times 8$), green ($4\times 4$), blue ($2\times 2$), and red ($1\times 1$). Higher-resolution patches focus on object centers for detail, while lower-resolution patches cover peripheral regions for efficiency. Mask predictions show accurate segmentation with optimized resource allocation.
3.4 Training
We train FLIP exclusively on META’s SA-1B Dataset Kirillov et al. (2023) with four model variants: FLIP-Tiny (0.51M parameters), FLIP-Small (2.3M parameters), FLIP-Middle (11.5M parameters), and FLIP-Large (96.6M parameters), performing approximately 8.5M, 7.4M, 5.8M and 2.7M updates respectively with a batch size of 256. These models are significantly more parameter-efficient than SAM variants (up to 641.1M parameters).
For data augmentation, we apply random perturbations to the 2D Gaussian derived from ground truth masks by shifting center coordinates, stretching or compressing $\sigma_{x}$ and $\sigma_{y}$, and introducing small rotations. The number of sampled tokens $N$ is varied around $\mu=512$ during training to improve generalization.
Sparse Mask Prediction.
Instead of dense mask prediction, we employ a computationally efficient sparse sampling strategy that focuses on boundary regions. Pixels are grouped into seven distance-based categories according to their L1 distance from mask boundaries: [0-2), [2-4), [4-8), [8-16), [16-32), [32-64), and [64$\infty$) pixels. For each sample, we select 2048 pixels distributed across these groups with equal sampling from inside and outside the mask to maintain class balance.
We implement adaptive sampling that dynamically redistributes pixels across distance groups based on prediction difficulty. Groups with lower IoU performance receive more samples in subsequent iterations, ensuring computational focus on challenging boundary regions.
The sparse mask loss reduces computational complexity from $O(HW)$ to $O(k)$ where $k\ll HW$:
$\mathcal{L}=\text{BCE}(\hat{M}_{\text{sparse}},M_{\text{sparse}})$ (1)
where $\hat{M}_{\text{sparse}}$ and $M_{\text{sparse}}$ represent predicted and ground truth values at sampled locations.
3.5 Inference
During inference, FLIP employs two strategies to efficiently generate segmentation masks while maintaining the sparse computation paradigm.
5-Sigma Bounding Box Prediction
Instead of predicting mask values for all pixels in the image, we exploit the Gaussian prior to restrict predictions to a 5-sigma bounding box around the object center. Given the Gaussian parameters $(\mu_{x},\mu_{y},\sigma_{x},\sigma_{y})$ from the input prompt, we compute a bounding box using $\sigma_{\text{iso}}=\max(\sigma_{x},\sigma_{y})$ and sample pixels only within the region $[\mu_{x}\pm 5\sigma_{\text{iso}},\mu_{y}\pm 5\sigma_{\text{iso}}]$. This reduces the number of mask queries from $O(HW)$ to $O(k^{2})$ where $k\propto\sigma_{\text{iso}}$, providing significant speedup for small objects while maintaining full accuracy for larger once.
Hierarchical Refinement
For further efficiency, we introduce a hierarchical inference strategy that progressively refines the mask prediction. Starting from a coarse resolution, we:
  1. 1.
    Predict initial mask values on the coarse grid
  2. 2.
    Identify uncertain regions where predictions fall inside $[\tau,1-\tau]$ (default $\tau=0.01$)
  3. 3.
    Upsample the mask by a factor $\alpha$ (default $\alpha=4$) using bilinear interpolation
  4. 4.
    Re-query only the uncertain pixels at the higher resolution
  5. 5.
    Repeat until reaching the target resolution
This approach dramatically reduces computational cost by focusing refinement on ambiguous boundary regions. The number of refinement steps $n$ is automatically determined as $n=\lfloor\log_{\alpha}(k_{\text{target}}/k_{\text{init}})\rfloor$, where $k_{\text{init}}$ and $k_{\text{target}}$ are the initial and target resolutions.
Both strategies can be combined, first using the 5-sigma bounding box to limit the spatial extent, then applying hierarchical refinement within that region. This enables real-time mask generation even for high-resolution images while maintaining maximum segmentation quality.
(a) SAM-H
(b) EfficientSAM-S
(c) MobileSAM
(d) FastSAM-x
(e) FLIP-Large
Figure 7: IoU (ObjaScale) heatmaps illustrating relative vs. absolute mask size. FLIP-Large retains strong accuracy even for small objects, provided there are enough pixels (e.g., $\geq 10\times 10$).
Figure 8: Comparison of mean IoU across six diverse datasets (Hypersim, KITTI-360, OpenImages, COCO, LVIS, ObjaScale) plotted against inference time (in milliseconds). FLIP variants consistently outperform SAM variants while being significantly faster.
Table 1: Comparison of Mean IoU (%) and Std IoU (%) across different datasets
Model Size (M) Time (ms) Hypersim KITTI-360 OpenImages COCO LVIS ObjaScale (ours) Average
Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%)
FastSAM-s 11.8 9.94 36.80 ± 34.62 37.71 ± 32.76 61.14 ± 32.10 46.51 ± 33.73 37.11 ± 35.19 48.20 ± 35.69 44.58 ± 34.02
FastSAM-x 72.2 24.32 34.36 ± 34.90 38.92 ± 34.06 69.31 ± 29.25 55.31 ± 34.55 43.24 ± 37.47 47.11 ± 36.41 48.04 ± 34.44
MobileSAM 10.13 21.15 68.03 ± 21.36 64.64 ± 18.87 82.20 ± 16.74 73.80 ± 17.48 71.59 ± 21.36 67.72 ± 25.03 71.33 ± 20.14
EfficientSAM-T 10.22 26.75 68.35 ± 24.52 60.82 ± 24.72 84.32 ± 15.08 75.93 ± 17.04 75.56 ± 19.95 68.74 ± 26.43 72.29 ± 21.29
EfficientSAM-S 26.41 47.98 69.65 ± 22.20 64.88 ± 19.35 85.99 ± 14.54 77.07 ± 16.81 75.18 ± 21.06 67.80 ± 26.76 73.43 ± 20.12
SAM-B 93.7 72.67 71.46 ± 20.88 62.38 ± 21.41 84.72 ± 15.38 76.07 ± 16.95 76.93 ± 19.17 71.38 ± 25.36 73.82 ± 19.86
SAM-L 312.3 148.78 72.13 ± 21.21 62.73 ± 20.31 86.94 ± 13.41 78.19 ± 16.05 77.93 ± 19.43 72.68 ± 25.22 75.10 ± 19.27
SAM-H 641.1 232.04 72.37 ± 21.65 62.47 ± 20.52 87.06 ± 13.53 78.41 ± 16.15 78.36 ± 19.50 73.76 ± 24.59 75.41 ± 19.32
FLIP-Tiny 0.51 9.82 73.17 ± 19.36 70.04 ± 15.27 83.83 ± 12.72 77.76 ± 13.55 76.88 ± 18.33 87.76 ± 14.93 78.24 ± 15.69
FLIP-Small 2.3 12.19 73.07 ± 19.80 69.91 ± 15.69 86.65 ± 11.27 79.23 ± 13.32 77.92 ± 18.34 88.96 ± 15.14 79.29 ± 15.59
FLIP-Middle 11.5 17.54 73.68 ± 20.13 69.86 ± 15.76 87.89 ± 10.08 80.16 ± 13.15 78.88 ± 18.31 89.08 ± 15.04 79.93 ± 15.41
FLIP-Large 96.6 38.65 73.66 ± 20.51 69.75 ± 16.03 89.33 ±   8.76 80.79 ± 13.06 79.42 ± 18.13 89.02 ± 15.21 80.33 ± 15.28
FLIPh-Tiny 0.51 9.05 71.36 ± 18.65 68.53 ± 15.08 83.03 ± 12.65 76.57 ± 13.25 75.41 ± 17.37 85.72 ± 14.52 76.77 ± 15.25
FLIPh-Small 2.3 10.94 71.93 ± 18.99 68.34 ± 15.57 85.73 ± 11.29 78.15 ± 13.04 76.71 ± 17.41 86.97 ± 14.68 77.97 ± 15.16
FLIPh-Middle 11.5 14.34 72.65 ± 19.21 68.28 ± 15.68 86.98 ± 10.01 78.98 ± 12.89 77.58 ± 17.25 86.95 ± 14.68 78.57 ± 14.95
FLIPh-Large 96.6 29.36 73.19 ± 19.56 68.05 ± 16.18 88.28 ±   8.96 79.67 ± 12.83 78.40 ± 17.10 87.36 ± 14.61 79.16 ± 14.87
Table 2: FLIP-Small ablation study
Ablation Params (M) Time (ms) Hypersim KITTI-360 OpenImages COCO LVIS ObjaScale (ours) Average
Mean ± Std Mean ± Std Mean ± Std Mean ± Std Mean ± Std Mean ± Std Mean ± Std
Baseline 2.3 12.19 73.07 ± 19.80 69.91 ± 15.69 86.65 ± 11.27 79.23 ± 13.32 77.92 ± 18.34 88.96 ± 15.14 79.29 ± 15.59
Baselineh 2.3 10.94 71.93 ± 18.99 68.34 ± 15.57 85.73 ± 11.29 78.15 ± 13.04 76.71 ± 17.41 86.97 ± 14.68 77.97 ± 15.16
InitialPosEmb 2.2 10.37 73.41 ± 19.73 68.96 ± 15.48 86.29 ± 11.54 79.22 ± 13.14 78.42 ± 17.98 89.05 ± 14.98 79.23 ± 15.47
InitialPosEmbh 2.2 8.98 72.03 ± 18.92 67.36 ± 15.27 85.45 ± 11.51 77.94 ± 12.98 76.92 ± 17.15 86.84 ± 14.63 77.76 ± 15.08
LaPE-SNorm 2.3 10.88 73.86 ± 19.37 69.92 ± 15.20 86.41 ± 11.40 79.29 ± 13.12 78.42 ± 18.00 89.12 ± 14.77 79.50 ± 15.31
LaPE-SNormh 2.3 8.79 71.03 ± 17.74 65.87 ± 13.78 80.20 ± 11.98 74.42 ± 12.71 74.34 ± 16.30 83.01 ± 14.61 74.81 ± 14.52
MaxPatch4x4 1.7 11.46 72.69 ± 19.71 69.43 ± 15.57 86.01 ± 11.55 78.73 ± 13.28 77.55 ± 18.30 88.48 ± 14.96 78.81 ± 15.56
MaxPatch4x4h 1.7 10.09 71.45 ± 18.97 67.87 ± 15.46 85.11 ± 11.47 77.61 ± 13.07 76.31 ± 17.43 86.55 ± 14.57 77.48 ± 15.16
MaxPatch64x64 3.0 12.54 73.45 ± 19.52 69.78 ± 15.93 86.67 ± 11.37 79.33 ± 13.15 78.39 ± 18.02 89.04 ± 14.89 79.44 ± 15.48
MaxPatch64x64h 3.0 11.10 72.19 ± 18.80 68.21 ± 15.80 85.73 ± 11.32 78.16 ± 12.91 77.04 ± 17.08 87.05 ± 14.41 78.06 ± 15.05
MinPatch4x4 2.2 11.45 69.39 ± 23.77 70.30 ± 15.60 86.42 ± 11.60 77.27 ± 16.62 72.08 ± 25.85 88.14 ± 17.28 77.27 ± 18.45
MinPatch4x4h 2.2 10.07 68.53 ± 22.58 68.82 ± 15.46 85.53 ± 11.60 76.31 ± 16.23 71.07 ± 24.74 86.05 ± 16.78 76.05 ± 17.90
NoDynamicSampling 2.3 12.17 71.03 ± 20.14 70.03 ± 15.46 87.47 ± 11.07 79.04 ± 13.65 76.40 ± 19.00 87.82 ± 15.95 78.63 ± 15.88
NoDynamicSamplingh 2.3 10.69 70.57 ± 19.46 68.54 ± 15.26 86.63 ± 11.07 78.29 ± 13.38 75.89 ± 18.11 86.66 ± 15.20 77.76 ± 15.41
Table 3: Segmentation Accuracy on small-scale objects ($<1$% image area)
Model Size (M) Time (ms) Hypersim KITTI-360 OpenImages COCO LVIS ObjaScale (ours) Average
Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%)
FastSAM-s 11.8 10.02 31.91 ± 32.85 32.55 ± 31.32 50.59 ± 32.66 38.50 ± 32.74 31.12 ± 33.01 45.91 ± 35.16 38.43 ± 32.96
FastSAM-x 72.2 24.37 24.77 ± 30.20 33.12 ± 33.34 56.58 ± 31.82 44.09 ± 34.05 34.96 ± 35.00 44.13 ± 35.62 39.61 ± 33.34
MobileSAM 10.13 21.83 65.67 ± 20.77 62.36 ± 18.41 76.76 ± 19.01 70.19 ± 17.36 68.46 ± 21.25 65.91 ± 24.86 68.22 ± 20.28
EfficientSAM-T 10.22 26.75 68.61 ± 20.18 63.11 ± 17.56 79.67 ± 16.78 72.91 ± 16.41 72.79 ± 19.80 67.11 ± 26.17 70.70 ± 19.48
EfficientSAM-S 26.41 47.98 66.74 ± 21.33 63.64 ± 18.01 80.82 ± 17.48 73.26 ± 16.78 71.64 ± 21.16 65.86 ± 26.63 70.33 ± 20.23
SAM-B 93.7 73.33 69.97 ± 19.58 63.05 ± 18.28 81.08 ± 16.87 74.48 ± 15.60 74.53 ± 19.11 69.79 ± 25.41 72.15 ± 19.14
SAM-L 312.3 149.45 69.76 ± 20.29 62.40 ± 18.44 82.76 ± 15.33 75.39 ± 15.59 74.75 ± 19.71 71.12 ± 25.32 72.70 ± 19.11
SAM-H 641.1 232.67 70.12 ± 20.66 62.02 ± 18.50 82.66 ± 15.81 75.62 ± 15.63 75.18 ± 19.83 72.26 ± 24.71 72.98 ± 19.19
FLIP-Tiny 0.51 8.58 71.04 ± 19.97 67.92 ± 15.59 81.53 ± 14.35 76.57 ± 13.93 75.18 ± 18.86 87.53 ± 15.16 76.63 ± 16.31
FLIP-Small 2.3 10.51 69.85 ± 20.17 67.53 ± 15.55 84.13 ± 12.70 77.18 ± 13.76 75.46 ± 18.89 88.75 ± 15.29 77.15 ± 16.06
FLIP-Middle 11.5 13.93 69.77 ± 20.39 66.69 ± 15.32 85.40 ± 11.21 77.69 ± 13.67 75.97 ± 18.98 88.86 ± 15.22 77.39 ± 15.80
FLIP-Large 96.6 29.89 69.18 ± 20.58 66.01 ± 15.39 86.85 ±   9.31 77.83 ± 13.47 76.12 ± 18.72 88.67 ± 15.51 77.44 ± 15.50
FLIPh-Tiny 0.51 8.21 68.94 ± 19.00 66.07 ± 15.14 80.26 ± 14.07 74.72 ± 13.31 73.45 ± 17.57 85.45 ± 14.72 74.81 ± 15.63
FLIPh-Small 2.3 10.01 68.71 ± 19.19 65.64 ± 15.17 82.58 ± 12.71 75.54 ± 13.10 74.09 ± 17.66 86.72 ± 14.79 75.55 ± 15.43
FLIPh-Middle 11.5 12.82 68.83 ± 19.34 64.78 ± 14.94 83.82 ± 11.08 75.92 ± 13.01 74.54 ± 17.55 86.71 ± 14.81 75.77 ± 15.12
FLIPh-Large 96.6 26.81 68.99 ± 19.63 63.94 ± 15.19 85.11 ±   9.47 76.21 ± 12.75 75.06 ± 17.40 86.99 ± 14.86 76.05 ± 14.88
(a) Fire Hydrant
(b) Apple
(c) Truck
(d) Airplane
Figure 9: Examples from our synthetic dataset. Objects from various categories—(a) Fire Hydrant, (b) Apple, (c) Truck, and (d) Airplane—are rendered with high-resolution HDRI Haven backgrounds. The dataset includes diverse objects and scene compositions, with varying object scales, resolutions, and viewing angles to challenge segmentation models. Best viewed zoomed in.
4 Results
We evaluate FLIP on five standard benchmarks: Hypersim, KITTI-360, OpenImages, COCO and LVIS Roberts et al. (2021); Liao et al. (2022); Kuznetsova et al. (2020); Lin et al. (2014); Gupta et al. (2019). Additionally, we evaluate FLIP on our newly constructed synthetic dataset ObjaScale, designed to stress-test scale invariance. Like SAM, FLIP was trained on the SA-1B dataset Kirillov et al. (2023), but was neither trained nor fine-tuned on the ObjaScale or any of the other evaluation datasets. Our experiments compare FLIP against state-of-the-art segmentation methods namely SAM Kirillov et al. (2023), EfficientSAM Xiong et al. (2023), MobileSAM Zhang et al. (2023), and FastSAM Zhao et al. (2023). Besides comparing the overall performance, we focus on handling small objects, achieving high IoU, and balancing parameter/runtime trade-offs.
4.1 Experimental Setup and Dataset Creation
We selected 68 diverse categories from Objaverse Deitke et al. (2023) and combined each with high-resolution HDRI Haven Zaal, Greg backgrounds in Blender, yielding 10,200 synthetic images, which form our ObjaScale dataset. Image resolution was randomized from $512$ to $8192$ pixels, while object dimensions were constrained to a maximum width/height of $256$ pixels, producing masks spanning minuscule ($<\!0.0001\%$) to large fractions of the image area (up to $25\%$) (Figure 9). For SAM variants, bounding-box prompts computed from ground-truth masks were used; FLIP employed 2D Gaussian prompts by design.
4.2 Overall Segmentation Performance
We denote hierarchical inference variants with the subscript “h” (e.g. FLIPh-Small); they share weights with their standard counterparts but perform progressive mask refinement (see subsection 3.5).
Table 1 presents comprehensive results across all datasets. FLIP consistently achieves superior performance compared to SAM variants while using substantially fewer parameters. FLIP-Large attains 80.33% mean IoU with 96.6M parameters, outperforming SAM-H’s 75.41% IoU that requires 641.1M parameters—a 6.6$\times$ parameter reduction with improved accuracy. Even our smallest model, FLIP-Tiny with only 0.51M parameters, achieves 78.24% mean IoU, surpassing all SAM variants including the 641M parameter SAM-H.
The performance advantage is particularly pronounced on our ObjaScale dataset, where FLIP models achieve 87–89% IoU compared to SAM-H’s 73.76%. This demonstrates FLIP’s superior scale invariance and ability to handle objects across diverse size ranges. On established benchmarks, FLIP maintains competitive performance: achieving 89.33% IoU on OpenImages (vs. SAM-H’s 87.06%) and 80.79% on COCO (vs. SAM-H’s 78.41%).
4.3 Parameter Efficiency and Performance
Figure 1 illustrates the parameter efficiency of FLIP variants. While maintaining superior segmentation quality, FLIP models operate with orders of magnitude fewer parameters than SAM variants. The efficiency gains translate to faster inference times (Figure 8), with FLIP-Large requiring only 38.65ms compared to SAM-H’s 232.04ms, representing a 6$\times$ speedup alongside improved accuracy (measured on a Nvidia RTX4090 graphics card).
Hierarchical Inference Trade-offs.
The hierarchical inference variants (FLIPh) demonstrate the effectiveness of our progressive refinement strategy, achieving comparable performance with reduced computational cost. Comparing standard inference with hierarchical variants reveals a general pattern: hierarchical inference achieves 98–99% of the original performance while providing $\approx$ 8–24% speed improvements. For instance, FLIPh-Large maintains 79.16% mean IoU while reducing inference time to 29.36ms. This trade-off is particularly favorable for real-time applications where slight accuracy reductions are acceptable for substantial computational savings.
4.4 Small Object Segmentation
Table 3 reveals FLIP’s capability for small object segmentation (objects $<$1% image area). FLIP-Large achieves 77.44% mean IoU on small objects compared to SAM-H’s 72.98%, demonstrating the effectiveness of our fovea-like patching mechanism. This advantage is most pronounced on ObjaScale, where FLIP-Tiny achieves 87.53% IoU on small objects compared to SAM-H’s 72.26%.
The heat-map analysis (Figure 7) offers clear evidence of FLIP’s scale-invariant behavior: whereas SAM variants—and FastSAM-x in particular—suffer sharp performance drops when the mask-to-image ratio falls below 0.01%, FLIP’s IoU remains stable until the object’s absolute footprint shrinks to roughly $10\times 10$ pixels. Because FLIP’s accuracy is governed by absolute pixel coverage rather than relative image area, it robustly segments small objects even in ultra-high-resolution scenes—a critical advantage for real-world applications.
4.5 Architectural Robustness and Ablation Analysis
Our ablation study (Table 2) reveals that FLIP is remarkably robust to architectural variations while certain design choices provide measurable improvements. The baseline FLIP-Small (79.29% mean IoU) serves as our reference point for evaluating architectural components.
Positional Embedding Strategies.
The comparison between baseline and “InitialPosEmb” (79.23% vs. 79.29%) shows that our layer-wise positional embedding injection provides marginal but consistent benefits over initial embedding only. However, the “LaPE-SNorm” variant, which applies shared normalization across positional embeddings, achieves the best performance at 79.50% mean IoU, while underperforming in the hierarchical inference setting.
Patch Resolution Configurations.
Testing different patch size ranges reveals interesting trade-offs. Restricting maximum patch size to 4$\times$4 (“MaxPatch4x4”) reduces parameters to 1.7M with only modest performance degradation (78.81% IoU), demonstrating FLIP’s efficiency even with limited resolution diversity. Conversely, extending to 64$\times$64 patches (“MaxPatch64x64”) slightly improves performance (79.44%) at the cost of increased parameters (3.0M), suggesting diminishing returns from very large patches.
Removing the smallest 1$\times$1 patches (“MinPatch4x4”) significantly impacts performance on LVIS (72.08% vs. 77.92%), highlighting the importance of fine-grained detail capture for complex segmentation tasks while maintaining strong performance on other datasets.
Sampling Strategy Impact.
The “NoDynamicSampling” ablation (78.63% vs. 79.29%) demonstrates that our adaptive pixel sampling strategy provides consistent benefits across datasets. Without dynamic sampling, each distance group uses the same number of pixels rather than adaptively redistributing based on prediction difficulty. The performance gap is particularly notable on Hypersim (71.03% vs. 73.07%), where focusing computational resources on challenging boundary regions better handles the diverse object shapes in simulated indoor scenes.
4.6 Cross-Dataset Generalization
FLIP demonstrates robust generalization across diverse evaluation benchmarks without dataset-specific fine-tuning. The performance gains are consistent across synthetic (Hypersim), automotive (KITTI-360), and large-scale natural image datasets (OpenImages, COCO, LVIS). Notably, FLIP-Tiny achieves 70.04% IoU on KITTI-360 compared to SAM-H’s 62.47%, despite using 1,257$\times$ fewer parameters. This 7.57 percentage point improvement highlights how FLIP’s fovea-like attention mechanism effectively handles automotive scenarios characterized by numerous small, distant objects.
The architectural ablations confirm that FLIP’s performance is remarkably robust to design variations. While removing the smallest patches (MinPatch4x4) impacts performance on complex datasets like LVIS, and the adaptive pixel sampling during training (NoDynamicSampling ablation) provides consistent improvements, the overall architecture maintains strong performance across different configurations. This robustness suggests that FLIP’s core fovea-like attention mechanism, rather than any single architectural choice, drives its superior accuracy, particularly for challenging small objects.
5 Conclusion
We have introduced FLIP, a novel vision model that fundamentally rethinks object segmentation through fovea-inspired input patching. By directly sampling multi-resolution patches centered on objects of interest, FLIP achieves state-of-the-art segmentation performance while using orders of magnitude fewer parameters than existing approaches.
Our key innovation lies in the scale-invariant patch sampling mechanism that adapts processing resolution to object characteristics, allocating computational resources where they matter most. This design enables FLIP-Large to achieve 80.33% mean IoU with only 96.6M parameters—outperforming SAM-H’s 75.41% IoU (641.1M parameters) while being 6$\times$ faster. Even our smallest model, FLIP-Tiny (0.51M parameters), surpasses all SAM variants with 78.24% mean IoU.
The introduction of ObjaScale reveals a critical limitation in current segmentation models: their inability to effectively handle very small objects in high-resolution scenes. FLIP addresses this gap through its fovea-like sampling, maintaining 87–89% IoU on ObjaScale where SAM variants achieve only 67–74%. This capability, combined with consistent performance improvements across Hypersim, KITTI-360, OpenImages, COCO, and LVIS, demonstrates FLIP’s potential for applications requiring precise segmentation of objects across extreme scale variations.
Looking forward, FLIP’s architectural efficiency and strong generalization open new possibilities for real-time object-centric vision systems. Its success suggests that biologically-inspired selective attention mechanisms can deliver both superior accuracy and computational efficiency, paving the way for more capable and sustainable computer vision applications in domains ranging from autonomous navigation to medical imaging.
6 Acknowledgement
This work received funding from the Deutsche Forschungs- gemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645 as well as from the Cyber Valley in Tübingen, CyVy-RF-2020-15. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Manuel Traub, and the Alexander von Humboldt Foundation for supporting Martin Butz.